OpenStudy (anonymous):

lim x-->0+ [1+(1/n)]^n

OpenStudy (anonymous):

This is a definition of \(e\), isn't it?

OpenStudy (anonymous):

\[ \lim_{x \rightarrow 0}(1+\frac{ 1 }{ n })^n\]

OpenStudy (anonymous):

You might wanna try using a logarithm...

OpenStudy (anonymous):

as in ln(1+1/n)^n = n ln (1+1/n)

OpenStudy (anonymous):

\[ 1+\frac{1}{n}=\frac{n}{n}+\frac{1}{n} = \frac{n+1}{n} \]

OpenStudy (anonymous):

is the n in front of the ln correct?

OpenStudy (anonymous):

Yes, but you also have to use the property for division inside logarithms.

OpenStudy (anonymous):

yes i see that now

OpenStudy (anonymous):

Then it is continuous. Remember the whole thing should be a power of \(e\).

OpenStudy (anonymous):

or maybe no continuous but rather... what's the term...

OpenStudy (anonymous):

but i can't have a n in the denominator with n approachhing zero

OpenStudy (anonymous):

One way I think about that part is to assume limit evaluates to some value L, then I take log of both sides.

OpenStudy (anonymous):

If you use the division property for logarithms, it is no longer divided.

OpenStudy (anonymous):

so its n[ln(n+1)-ln(n)]?

OpenStudy (anonymous):

yes, no?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

well what do we do with that n on the outside? it's negating the entire equation

OpenStudy (anonymous):

Nothing...

OpenStudy (anonymous):

that would give us ln y = 0

OpenStudy (anonymous):

I'm much more concerned with \(\ln(n), n\to 0\) because

OpenStudy (anonymous):

It doesn't exist going from the left.

OpenStudy (anonymous):

i don't understand, is that what that plus is for?

OpenStudy (anonymous):

ok i'm at \[lny=\lim_{x \rightarrow 0+} n]\ln(n+1)-\ln(n)]\]

OpenStudy (anonymous):

am i anywhere near the right track?

OpenStudy (campbell_st):

actually there is no limit... because if x===> 0 in an expression without x's you don't need to do anything just an observation...

OpenStudy (anonymous):

i meant n->0+

OpenStudy (anonymous):

now can you help?